OpenStudy (anonymous):

How many ppm (parts per million) are there in one percent (1%)?

OpenStudy (anonymous):

this is the same as asking, how many millionths are there in one one-hundredth, so we divide one one-hundredth by one one-millionth:\[\frac{ \frac{ 1 }{ 100 }}{ \frac{ 1 }{ 1000000 } } = \frac{ 1000000 }{ 100 } = 10000\]